Find the center of mass of four 2 kg bodies placed at each corner of a square of side 1.5 m
1
Expert's answer
2020-12-14T12:17:20-0500

Let the left bottom body be at the origin.


"4mX=m(0+0+1.5+1.5)\\X=0.75\ m
\\4mY=m(0+0+1.5+1.5)\\Y=0.75\ m\\"

 COM is at center of square.